OpenStudy (anonymous):

Could anyone help me solve this? 9^-x times 12^x

OpenStudy (anonymous):

\[9^{-x} \times 12^{x} = \frac{ 12^x }{ 9^{x} } = (\frac{ 12 }{ 9 })^x\]

OpenStudy (anonymous):

12/9 can be simplified further into 4/3

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Do you know how to do \[\sqrt{3} + \sqrt[2]{2} \div \sqrt{3} - \sqrt[2]{2}\]